# Non-linear conductances of Galton–Watson trees and application to the (near) critical random cluster model

### v1

@@ Appendix B Branching processes with (truncated) heavy tails
$\begin{split}\mathbf{P}\big{(}W_{\ell+1}\geqslant c_{\ell+1}x\big{)}&\leqslant\mathbf{P}\big{(}W_{\ell}\geqslant c_{\ell}x\big{)}+cm^{-\frac{1}{2}(\alpha-1)\ell}L(x)x^{-\alpha}+cm^{-\alpha\ell/2}\mathbf{P}\big{(}W_{\ell}\geqslant\tfrac{1}{2}c_{\ell}x\big{)}\,,\end{split}$ (B.4)
@@
for some slowly varying function $\hat{L}(\cdot)$. Now, if we take $\lambda\leqslant c_{\alpha}\frac{\log t}{t}$ with a constant $c_{\alpha}:=\frac{1}{2}(\alpha\wedge 2-1)>0$, we get that $\varepsilon_{\lambda}\leqslant c_{\alpha}\log t\hat{L}(t)t^{-2c_{\alpha}}t^{c_{\alpha}}$, so in particular $\tilde{\varepsilon}_{\lambda}\leqslant 1$, provided that $t$ is large enough.
Let us fix $\lambda=\delta\frac{\log t}{t}$ with $\delta:=c_{\alpha}\prod_{i\geqslant 1}(1+\tilde{m}^{-i})^{-1}$, and let us define $(\lambda_{k})_{1\leqslant k\leqslant\ell}$ by setting $\lambda_{1}=\lambda$ and $\lambda_{k+1}:=(1+\tilde{m}^{k-\ell})\lambda_{k}$ for $11\leqslant k\leqslant\ell-$ 1. Notice that $\lambda_{k}\leqslant\lambda_{\ell}\leqslant c_{\alpha}\frac{\log t}{t}$; in particular $\varepsilon_{\lambda_{k}}\leqslant 1$ for all $k\in\{1,\ldots,\ell\}$. Applying (B.8) iteratively, we then get that
@@
Acknowledgement.
We would like to thank Remco van der Hofstad, Piet Lammers, Arnaud Le Ny, Yueyun Hu and Raphaël Rossignol for several interesting and enlightening discussions on the subject.

### v2

@@ Appendix B Branching processes with (truncated) heavy tails
[…]
@@
[…]
Let us fix $\lambda=\delta\frac{\log t}{t}$ with $\delta:=c_{\alpha}\prod_{i\geqslant 1}(1+\tilde{m}^{-i})^{-1}$, and let us define $(\lambda_{k})_{1\leqslant k\leqslant\ell}$ by setting $\lambda_{1}=\lambda$ and $\lambda_{k+1}:=(1+\tilde{m}^{k-\ell})\lambda_{k}$ for $1\leqslant k\leqslant\ell-1$. Notice that $\lambda_{k}\leqslant\lambda_{\ell}\leqslant c_{\alpha}\frac{\log t}{t}$; in particular $\varepsilon_{\lambda_{k}}\leqslant 1$ for all $k\in\{1,\ldots,\ell\}$. Applying (B.8) iteratively, we then get that
@@
Acknowledgement
We would like to thank Remco van der Hofstad, Piet Lammers, Arnaud Le Ny, Yueyun Hu and Raphaël Rossignol for several interesting and enlightening discussions on the subject; we also thank Christina Goldschmidt and Markus Heydenreich for communicating to us that they obtained the (same) constants in Theorem 2.6 in an unpublished work. Finally, we are grateful for the insightful comments of two anonymous referees, who helped us improve the presentation of the paper. Q. Berger acknowledges the support of grant ANR-22-CE40-0012.